= Problem 2: Let I(f) = S• f (x)dx. We are interested in approximating this integral within a certain error tolerance. Firstclass : numerical analysis

I wish if it was written in block letter

Sorry I can't read cursive
Answer #1

WE HAVE 11 ERROR = \Eglo = | I (f) - Mo () | T 1 h² fic CE(a,b) lent (b-a) fucc) 24 n ㅗ 24 3 HERE I = Lex? dx SO a=0 b = 3 9f(x) = – 2x ex? +6 f(x) = + 4x?e *? 2xe-x = (4x²-2) e-x² 4 XE (03) 3 Take g (x)= (4x²-2) e-x² g(x) = (4x2_2) e-X² (-2x) te3 1 FOR a> > g(x) 1 eves cave) 1 - ve 1 sog(x) gcxCHANGES FROM +ve To -ve X = 3 CORRESPONDS TO MAXIMA MAX VALUE95 st 3 c/4